A firm demand function for its output is p=400-0.5q. To maximize sales revenue,what price should be charged


1
Expert's answer
2021-11-18T10:19:46-0500

Given the general form of Demand Function: P = f(Q),

Then the general form of Inverse Demand Function is: O = f-1(P)

 Demand Curve: "P = 400 - 0.5 Q"

Driving Inverse Demand Curve:

"0.5 Q = 400 - P\\

Q =\frac{ (400 - P ) }{ 0.5}\\

Q = 400 / 0.5 - P / 0.5"

Inverse Demand Curve: Q = 800 - 2P


We can see that at Price $200, the total revenue is maximized at TR = $80,000

=> To maximize sales revenue, firm should charge price of $200